OpenStudy (newwar):

What is the value of f(−1)f(−1) when f(x)=2x+2f(x)=2x+2 ? Enter your answer in the box. f(−1)=f(−1)=

OpenStudy (518nad):

multiply f(x)*f(x)

OpenStudy (518nad):

f(x)*f(x) = (f(x))^2=(2x+2)^2 (f(-1))^2=(2(-1)+2)^2 =0

OpenStudy (518nad):

or just do this f(-1)=2(-1)+2 =0 so f(-1)=0 then f(-1)*f(-1) = 0 * 0 = 0

OpenStudy (newwar):

so the answer is 0 @518nad